OpenStudy (anonymous):

A square is inscribed in a circle. How fast is the area of the square changing when the area of the circle is increasing at the rate of 1 in^2/min?

OpenStudy (p0sitr0n):

A circle = pi r^2 dA = + pi d (r^2) / dt dA/dt = + pi 2r dr /dt (dA/dt ) / 2pi r = dr / dt r^2=2c^2 r^2/2=c^2 = A square r dr /dt = dA square /dt substitute dr (dA/ dt) / 2pi = dA square / dt where dA /dt is your rate of change of circle area